OpenStudy (anonymous):

@jim_thompson5910 What is the exact value of the expression 8the square root of 15. + 5the square root of 5. + the square root of 15.? 14 sqrt 35 ?

jimthompson5910 (jim_thompson5910):

combine like terms

jimthompson5910 (jim_thompson5910):

this is NOT a valid move \[\large \sqrt{15}+\sqrt{15} = \sqrt{30}\] that equation above is NOT true

jimthompson5910 (jim_thompson5910):

so, \[\large \sqrt{15}+\sqrt{15} \neq \sqrt{30}\]

OpenStudy (anonymous):

Okay.

jimthompson5910 (jim_thompson5910):

it might help to replace every sqrt(15) with x and sqrt(5) with y

jimthompson5910 (jim_thompson5910):

that would turn \[\large 8\sqrt{15} + 5\sqrt{5} + \sqrt{15}\] into \[\large 8x + 5y + x\]

OpenStudy (anonymous):

13xy+x?

jimthompson5910 (jim_thompson5910):

8x+5y cannot be combined to 13xy

jimthompson5910 (jim_thompson5910):

x and y are unlike terms

jimthompson5910 (jim_thompson5910):

but 8x and x are like terms

OpenStudy (anonymous):

9x

jimthompson5910 (jim_thompson5910):

so you now have 9x + y

jimthompson5910 (jim_thompson5910):

that turns back into \[\large 9\sqrt{15} + \sqrt{5}\]

jimthompson5910 (jim_thompson5910):

so, \[\large 8\sqrt{15} + 5\sqrt{5} + \sqrt{15} = 9\sqrt{15} + \sqrt{5}\]

OpenStudy (anonymous):

Oh, thanks again.

jimthompson5910 (jim_thompson5910):

yw